OpenStudy (anonymous):

Help. how do you solve the area of rugby field = 8160. How do I find the dimensions of the field if the length of the rugby field is 52 meters longer than it's width

OpenStudy (gw2011):

Area=(length)(width) Area=8160 Width=w Length=52+w 8160=(w)(52+w) 8160=52w+w^2 Rearranging this equation you get: w^2+52w=8160 w^2+52w-8160=0 Factoring this equation you get: (w+120)(w-68)=0 Solve for (w-68) because this will be a positive number w-68=0 w=68 l=68+52=120 Therefore, Width=68 meters Length=120 meters You can verify this by the following: Area=( Length)(Width) Area=(120)(68)=8160 meters